FIND THREE CONSECUTIVE ODD INTEGERS SUCH THAT THREE TIMES THE SECOND MINUS THE THIRD IS 17 MORE THAN THE FIRST.
Answer by sudhanshu_kmr(1152)


x, x+2 and x+4
3(x+2) - (x+4) = 17 + x
=> 3x + 6 -x -4 = 17 +x
=> x = 15
so, numbers are 15, 17 and 19.